OpenStudy (camerondoherty):

A system of equations is shown below: 8x + 5y = 9 3x + 2y = 4 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (dan815):

8x + 5y = 9 3x + 2y = 4 ---------- (8+3)x+(5+2)y=9+4 11x+7y=13

OpenStudy (camerondoherty):

???

OpenStudy (camerondoherty):

just explain to me what u did xD

OpenStudy (dan815):

i added the first and 2nd equation together

OpenStudy (camerondoherty):

o ok

OpenStudy (dan815):

because umm okay.. an* equation has 2 sides left 1= right 1 left 2 = right 2 if i add left 2 to left 1.... and if i add right 2 to right 1... that still the same as adding left 2 to both sides since left 2 = right 2

OpenStudy (dan815):

now you have to make another equation by taking a multiple of one of those equation , for example 2*equation2= 2*(3x+2y)=2*4 6x+4y=8

OpenStudy (camerondoherty):

ooooooo ok

OpenStudy (camerondoherty):

thanx a bunch :)

OpenStudy (camerondoherty):

can you re-explain tht part again?

OpenStudy (camerondoherty):

i dont really get it...

OpenStudy (camerondoherty):

@dan815

OpenStudy (dan815):

the addition of equations or the multiple of an equation

OpenStudy (camerondoherty):

o.o

OpenStudy (camerondoherty):

english?

OpenStudy (dan815):

oh hey I just reread your equation and I think they just want you to replace one of the equations

OpenStudy (dan815):

okay so lets take the sum of one equations and a multiple of the other one

OpenStudy (dan815):

okay sooo loook pay attention!!

OpenStudy (camerondoherty):

ok

OpenStudy (dan815):

equation 1 equation 2 Lets say our new equation 3=equation1+2*equation2

OpenStudy (camerondoherty):

what does * mean?

OpenStudy (dan815):

8x + 5y = 9 (1) equation 3x + 2y = 4 (2) equation -------------- (3rd equation)=(1st equation)+2*(2nd equation)

OpenStudy (dan815):

* means multiplication

OpenStudy (camerondoherty):

just put times or (space)x(space)

OpenStudy (dan815):

New equation = 8x + 5y = 9 +2(3x+2y)=4 ------------ 14x+9y=17 <--------- new equation -----------------

OpenStudy (dan815):